Please Answer all these. Those questions are from New Perspectives Microsoft Office 365 & Excel 2016 .

Need answers from Excel concern :

1. What formula would you enter to add the values in cells C1, C2, and C3?
What function would you enter to achieve the same result?
2. What formula would you enter to count how many numeric values are in the
range D21:D72?
3. If you insert cells into the range C1:D10, shifting the cells to the right, what is
the new location of the data that was previously in cell F4?
4. Cell E11 contains the formula =SUM(D1:D20). How does this formula change
if a new row is inserted above row 5?
5. Describe four ways of viewing the content of a workbook in Excel.
6. How are page breaks indicated in Page Break Preview?
7. What orientation would you use to make the printed page wider than it is tall?
8. How do you display the formulas used in a worksheet instead of the formula
results?
Answer #1

1. Use formula: =C1+C2+C3

or Use sum function: =SUM(C1:C3)

2. Use formula: =COUNT(D21:D72)

3. The new location of data is cell G4.

4. If new row inserted above row 5 then formula will become =SUM(D1:D21).

5. Four ways of viewing content of workbook in Excel are:

a. Normal View: This view is used to show content of workbook.

b. Page Layout : This view shows how the workbook will look when get printed.

c. Page Break Preview: This view is used to show locations of various page breaks in the workbook.

d. Formula View : This view is used to show the formulas used in the workbook.

6. Page breaks indicated in page break preview by a blue dotted border.

7. Landscape orientation would be used for this purpose.

8. For this purpose we can hold the Ctrl + , it will show the formulas or we can go to Formulas in the ribbon and click on 'show formulas' button.